OpenStudy (anonymous):

integerate (1+sin2x)^1/2 dx

hartnn (hartnn):

write 1 as \(1=\sin^2x+\cos^2x\) and \(\sin 2x = 2\sin x \cos x\)

hartnn (hartnn):

and get surprised that you'll get a perfect square! :)

OpenStudy (anonymous):

yeah i got it. thanks again.

hartnn (hartnn):

welcome ^^